OpenStudy (anonymous):

write the slope intercept equation for the line that passes through (-4,-8) and (-3,-2).

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (aroub):

-2+8/-3+4 6/1 6 y=mx+b , m=6 , (-4,-8) -8=6(-4)+b -8=-24+b 24-8=b 16=b so the equation is y=6x-16